x-{3x-[2x-(8-x)+2]+3}-6=0
Answer by jim_thompson5910(35256)

x-{3x-[2x-(8-x)+2]+3}-6=0 ... Start with the given equation.


x-{3x-[2x-8+x+2]+3}-6=0 ... Distribute the inner most parenthesis


x-{3x-[3x-6]+3}-6=0 ... Combine like terms.


x-{3x-3x+6+3}-6=0 ... Distribute


x-9-6=0 ... Combine like terms.


x-15=0 ... Combine like terms again


x=15 ... Add 15 to both sides.


So the answer is x=15